A multiple choice test consists of 8 question; each with 4 possible answers but only one of which is correct. In how many ways can a student answer the 8 question and
a. Get them all correct?
b. Get them all wrong?
c. Fet atleast one correct answer?

Answer by ewatrrr(24785)

p(correct) = 1/4 and p(wrong) = 3/4
1.P(all correct) = (1/4)^8
2.P(all wrong) = (3/4)^8
3. P(at least one correct) = 1 - (3/4)^8